UT 2018 09-23 D/E Split Night w/SOUL


Observers: B. Rothberg (Tucson)
OSA: S. Allanson

ISp: E. Solheid

Mtn Manager: P. Hartley

OPEN

UT 07:17 - extra time for SOUL work

UT 08:18 - handing over to D/E time with 2-4" seeing

UT 08:22 - reconfiguring for MODS-1

UT 08:27 - LUCIs put in safe config

UT 08:38 - sending preset to Feige 110

UT 08:49 - taking spectra

UT 08:51 - Red IMCS lock failed - probably because we are at El=39, retry worked

UT 09:05 - recovering from primary mirror panic

UT 09:10 - pointing and collimating near next target HV Cet

UT 09:16 - we hit 5" on the DIMM - having trouble finding the centroid and flux on an R=14
UT 09:20 - changed to 1.2" slit for HV Cet given the variable seeing.

UT 09:30 - starting spectroscopy of HV Cet 1.2" slit

-the bright mottling, fringing? seen in the acquisition is visible in the bottom half of the slit.

UT 10:34 - seeing has dramatically improved in short order <0.9" now

HV Cet done mods1b 0005-007, mods1r 0010-012

UT 10:38 - moving to RQQ Mrk 1014

UT 10:40 - and back to 1.5-2" in the guider

UT 10:52 - need to nudge left -0.123" DETX

UT 10:55 - starting spectroscopy

UT 11:29 - taking a 2nd 30min exposure Dual

UT 12:02 - taking a 15min exposure Dual

UT 12:19 - finished, enclosure closing - will take 1.2" slitflats

UT 12:29 - starting cals 1.2" slitflats mods1b.0011-0016 , mods1r.0020-0025

-- %USERSIG{BarryRothberg - 2018-09-23}%
Topic revision: r2 - 23 Sep 2018, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback